PartnerObserving Log for 2007 Mar 18/19

Observers: Rick Pogge, Eichii Egami
Support Scientist: none
TO: Aaron Ceranski

Lots of high clouds around during the afternoon, mostly moving through, but plenty of junk off to the WSW on the satellite to keep us guessing tonight. The forecast calls for mostly clear tonight and cooler, but the clouds coming up out of Mexico (appearing to pickup moisture while passing over the Gulf of California) are not giving me warm fuzzies about our chances right now.

All times below are UTC 2007 March 19

0130 - Taking evening bias and dark frames. Weather is cloudy with scant patches of blue between a series of moving bands of thick high clouds.

0230 - Opened, going to the GRB07311 field to setup mirror, but clouds are iffy. T=6C, Wind 8m/s out of the NE, Humidity 27%. Adjusting mirror with Rfastextra. taking a while to converge.

0250 - 5s r shot to ID field and make sure we aren't saturating the reference star. Good for 20s. 0.8" seeing, but images are out of round

0255 - GRB/GRB07331, 5x20s calibration snaps
  • Image 1: 025611 - 0.73" FWHM, peak in Refstar A is 30000ADU
  • Image 2: 025708 - 0.84" FHWM, peak=21700
  • Image 3: 025808 - 0.9" FWHM, peak=21000
  • Image 4: 025905 - 0.75" FWHM, peak=16500
  • Image 5: 030011 - 0.75" FWHM, peak=5300

0310 - Losing the field behind passing clouds. Walking outside, it is hard to see Orion proper. Unless things change, we won't be able to get a deeper image of this field. At least this gives the GRB team a way to scale the deep images from the previous night to data taken by observers at other telescopes.

0322 - shot a couple more snaps through holes in the clouds around Orion. Seeing now 1-1.2". Going to be hard to decide on targets given clouds and seeing. Will execute the "poor-seeing" (>1") programs for now.

0334 - Skillman/NGC2366 (seeing is 1.0-1.2") - tweaking primary with Bfastextra, started NGC2366_U1 at 0341
  • Image 1: 034223 - has a guide star (looks a elongated at times) - 1.1" seeing, stars not as round as they could be
  • Image 2: 034809 - gsview showing garbage, but appears to be guiding - 0.9" seeing and round
  • Image 3: 035349 - 1.0" seeing and nice and round.
  • Image 4: 035934 - gs lock failed about half-way through 300s exposure - 1.1" seeing, images a little elongated as a result of ~180s of open-loop tracking

GUI froze up during first minute of Image 3 above all by itself, and didn't come back until about 50s into the 4th image (the change views trick didn't work). The archive, even for tech chips, was about 8 minutes behind at one point, then started delivering them in essentially random order, recent ones interspersed with older ones. Image 2 appeared part way through the integration of Image 4, while Image 3 didn't appear until just before Image 4 appeared. Not sure this info helps anyone debug the system, but there you have it.

0404 - Tweaking primary (some delays) - starting NGC2366_U2 at 0409 (delay of mirror reset and instrument config)
  • Image1: 041052 - 1.15" seeing, images reasonably round
  • Image 2: 041632 - 1.15", round, but images are fainter by ~2x - clouds coming in on us
  • Image 3: 042211 - gs failed 180s in, image getting clobbered by clouds
  • Image 4: 042756 - aborted, target is behind clouds. wide-field MAT shot showed no stars (about 3-deg field).

0429 - Had to abort the OB because of clouds. Satellite shows a big patch coming over us now, but there isn't as much stuff coming towards us from the WSW as before.

0447 - Stars starting to appear on the MAT, checking LBT focus. Starting NGC2366_U3, as conditions appear to be improving.
  • Image1: 045341 - guide star puffing up, but transparency better - images have 1.4" images, but round - flux about half what it was at start of previous OB.

Aborting the rest of the sequence due to clouds and poor seeing. Also getting past its airmass limit for U, so moving on to other targets.

0500 - PartnerObserving Dwarfs/mpobj3, which is tolerant of >1" seeing. Won't know if it is clear until we get there... taking 5s G snapshot to see. Tweaking primary with Gfastextra, Took a couple more iterations than usual

0515 - Starting with OB mpobj3_1_g
  • Image1: 051617 - seeing variable (judged from gs), but looks like 1.1" on images, reasonably round. Background a little high (9000ADU)
  • Image2: 052159 - seeing still variable, 1.2" on the images. background a little lower, more like 7200ADU now
  • Image3: 052745 - losing guide star every now and then, background up (10000ADU), and about 2 mag fainter on our stars
  • Image4: 053344 - clouds coming in, image clobbered - all stars are doubled, definitely an unhappy guider

0545 - MAT images showed steadily deteriorating transparency through this sequence, even though we started out pretty good during the mirror tweak-up. It's going to be that kind of night... We were between two streams of clouds earlier, but then those streams converged right over us. It may take a while to pass.

0649 - Getting fairly windy and very cloudy, we're closing the doors for now. Could be a while before we will open again, very overcast.

0750 - Only a little improvement - you can see a few bright stars, but not enough for a whole constellation. Windy and colder, too.

0850 - Little change, clouds covering most of the sky, where you can see stars there is maybe 2mag of extinction.

0915 - Sky looking a little bit better to the South and East, but still lots of junk up there. Going to try to see if we can resume observing. Pointing to the Dwarf/mpobj1 field, and running Gfastextra to tweak the focus/collimation.

0922 - Dwarfs/mpobj1, starting with OB mpobj1_1_g.ob
  • Image1: 092355 - gs lost lock - aborted, checked the technical chips, there are stars, so restarted 092727 - found a star this time and locked, lost it, then recovered, then lost it, then recovered, and so on. The seeing is not especially bad or variable, and the gs shown looks pretty good. Seeing is ~1.25" and reasonably round images.
  • Image2: 093312 - better guiding - 1.1" and round. Counts same as in Image1, so we're doing better with transparency
  • Image3: 093905 - no gs problems, 1.1", very round. fluxes still stable
  • Image4: 094503 - no problems, seeing 1.1"+, slightly rattier profiles, fluxes stable

0950 - Tweaking primary, then mpobj1_1_r.ob,
  • Image1: 095724 - 1.2" images, reasonably round
  • Image2: 100311 - 1.2" images, roundish stars
  • Image3: 100856 - satellite streak across Chip 1 & 2, 1.1+" images, very round, fluxes stable
  • Image4: 101454 - lost guider lock early in exposure - 1.2" and slightly elongated.

1020 - Tweaking primary, then mpobj1_2_g.ob
  • Image1: 102605 - 1.3" images, but very round
  • Image2: 103152 - gs lost lock - images degraded and lop-sided, typical width ~1.5"
  • Image3: 103742 - gs lost lock - same ugly images as 103152 -
  • Image4: 104338 - gs held this time, high background (clouds) - 1.1+" seeing, very round

1049 - Tweaking primary, then mpobj1_2_r.ob
  • Image1: 105634 - 1.2", a little out of round
  • Image2: 110221 - 1.1" and very round
  • Image3: 110805 - 1" and very round
  • Image4: 111359 - 1" and very round

1127 - GUI locked up trying to download an OB. Server appears to have crashed. Attempt to restart the GUI failed. Killed and restarted the server, then restarted the GUI. Recovered at 1137.

1138 - Tried to load an OB, and it blinked the GUI once and shut down the IIF link, then the instrument. All are "off" on the Power Control screen. Again. Restarted housekeeping, then restarted instrument systems, camera, and IIF link. Back at 1148. Took 5s g snapshot, seems to be working again, but we burned up the last 1/2 hour of the night

1250 - Trying to setup skyflats, but too many clouds around, showing up as moving structures on the images. No point in trying to do sky flats.

Night Summary

Clouds over most of the sky most of the night. A few windows where we could observe targets. Seeing started as good as ~0.7" but quickly degraded to around 1.0-1.2" or worse. Very little observing.

Programs Observed:
  • GRB/GRB07311 - got brief calibration snapshots (5x20s) in deteriorating conditions, but sufficient to do the relative calibration.
  • Skillman/NGC2366 - only completed U1 and part of U2 before clouds stopped observing
  • Dwarfs/mpobj3 - only got part of the first OB (mpobj3_1_g) before clouds stopped observing
  • Dwarfs/mpobj1 - completed most of the 4 OBs, but some images were affected by loss of guider lock in the 3rd set (mpobj3_2_g)

Forecast is looking better for Monday night.

-- RichardPogge - 18 Mar 2007
Topic revision: r9 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback